OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1 A (-3, -1, -5) B (-3, 5, -2) C (2, -4, -3) D (4, 4, -1)

OpenStudy (kenljw):

-1 (x + 3y + 2z = -16) 2x - y + 2z = -15 x -4y = 1 x = 1 + 4y 2x - 2y - z = 1 z = 2x -2y -1 z = 2 (1 +4y) -2y -1 z = 6y + 1 : y < -1/6 z negative can't be B , C, D y = -1 gives z = -5 checks y = -1 gives x = -3 Therefore A is answer